immacoolperson:

what is -3/7 divided by 9/4?

Max1oof:

I think it Is -12/63, but I am not 100% certain

immacoolperson:

thanks, that's what I got too.

dude:

Yeah multiply by the reciprocal \(\dfrac{-3}{7}\times \dfrac{4}{9}\)

Max1oof:

okay well im gonna say if two people got it, its probably right and the 3 isnt negative, the entire fraction is im thinking, dude

dude:

Ye negative on the numerator or denominator still makes the whole fraction negative, it doesnt matter in this case

Max1oof:

yeah I really suck at that whole thing lmao so ur gonna have to bear with me on that : |

immacoolperson:

now I need help with this one: -12 divided by 0.03. sorry, super bad at math

Max1oof:

I am too lmao

dude:

Do you know what 0.03 is as a fraction?

immacoolperson:

I think its 3/100 ?

dude:

Yes good So it's asking what -12 divided by \(\dfrac{3}{100}\) is You first find the reciprocal of the second fraction (Aka just flip the fraction) Because you have the reciprocal, just multiply both \(-12\times\dfrac{100}{3}\)

immacoolperson:

1200/30 ?

dude:

Divided by 3, not 30

immacoolperson:

ohhh ok

immacoolperson:

400 doesn't seem right....

dude:

Whoops meant \(-12 ~\div~0.03=-400\)

immacoolperson:

ok. thank you